OpenStudy (anonymous):

Tarzan, who weighs 875N, swings from a vine through the jungle. How much work is done by the tension in the vine as he drops through a vertical distance of 4m?

OpenStudy (anonymous):

how to draw and describe this?

OpenStudy (shamim):

W=force*distance

OpenStudy (shamim):

Given force=875N Distance=4m W=?

OpenStudy (anonymous):

The answear is zero, I need to know why!

OpenStudy (shamim):

W=875*4=?

OpenStudy (anonymous):

No there is none, there is zero work done by the tenision in the vine! but why?

OpenStudy (shamim):

Ok

OpenStudy (shamim):

When tarzan will go downward, the work by tension is negative -875*4j And when the tarzan go upward , work done by tarzan is positive +875*4j

OpenStudy (shamim):

Total work done by tension is=-875*4+875*4=0

OpenStudy (shamim):

Response plz

OpenStudy (anonymous):

Yea that could be it :)

OpenStudy (shamim):

Medal?!!!!